A courier service company wishes to estimate the proportion of people in various states that will use its services. Suppose the true proportion is 0.06. If 305 are sampled, what is the probability that the sample proportion will be greater than 0.08? Round your answer to four decimal places.
Answer #1
here population proportion=     μp= 0.0600
sample size       =n= 305
std error of proportion=σp=√(p*(1-p)/n)= 0.0136

probability that the sample proportion will be greater than 0.08 :

probability =P(X>0.08)=P(Z>(0.08-0.06)/0.014)=P(Z>1.47)=1-P(Z<1.47)=1-0.9292=0.0708